OpenStudy (anonymous):

PLZ HELP USing an appropriate substitution, solve \[(x^2 tan y)y'=x^6 sin (3x)+3x ln (sec y)\]

OpenStudy (dan815):

-.-

OpenStudy (anonymous):

I need help

OpenStudy (anonymous):

Try \(t=\ln \sec y\), then \(t'=\dfrac{\sec y\tan y}{\sec y}y'\), or \(t'=\tan y~y'\). \[\begin{align*} x^2\tan y~y'&=x^6\sin3x+3x\ln\sec y\\\\ x^2t'&=x^6\sin3x+3xt\\\\ t'-\frac{3}{x}t&=x^4\sin3x \end{align*}\] which is linear in \(t\).